OpenStudy (anonymous):

Derrick charges a flat rate of $10 for babysitting, plus an additional $4 per hour. If "h" represents the number of hours he babysits, which algebraic expression represents the total charge?

sam (.sam.):

Total charge IS immediate(flat) of $10 PLUS 4 per hour C = $10 + 4/hour C= 10 +4h C= 4h +10